A copper wire is stretched to make it 0.2% longer. What is the percentage change in its resistance?


1
Expert's answer
2022-01-17T09:59:49-0500

"R\sim \frac lS,"

"l_1=1.02l,"

"S_1=0.98 S,"

"R_1\sim\frac{1.02l}{0.98S}=1.04\frac lS,"

"R_1=1.04R," 4%.